OpenStudy (anonymous):

Help solving 2x^3/4=54?

OpenStudy (anonymous):

2x^3/4=54 (1/2)(2x^3/4)=(1/2)(54) x^(3/4)=27 (x^(3/4))^4 = 27^4 x^3=27^4 (x^3)^(1/3)= (27^4)^(1/3) x=81

OpenStudy (anonymous):

Thank you! Very helpful :)